# Tight MIP Formulations for Optimal Operation and Investment of Storage Including Reserves

### v2

@@ II-C Modeling Storage Investment Problems Including Reserves
We can extend the BOR-MIP model, as presented in the previous section, to model investment decisions. The most general way to model the investment of storage is the Basic Investment and operation incl. Reserves MIP model (BIR-MIP model) [21]. The operational problem is similar to that in the BOR-MIP model, but it contains additional investment variables $\overline{e}$, $\overline{c}$, and $\overline{d}$. The capacity limits of the storage unit, which now depend on the investment variables, are tracked in (3a) and (3b), where parameter $\overline{E}_{0}$ is the initially installed maximum storage capacity. Parameter $\theta$ represents the size of the minimum storage capacity as a fraction of the maximum. The upper bounds for charging and discharging and the reserves, also dependent on the investment variables, are imposed in (3c)-(3f), where parameters $\overline{P^{\mathrm{C}}_{0}}$ and $\overline{P^{\mathrm{D}}_{0}}$ are the initially installed charge and discharge capacities, and $C$ and $D$ are the maximum amount of invested charge/discharge capacity. Bounds on the maximum amount of invested capacity are imposed in (3g)-(3i), where $C$, $D$, and $E$ are typically quite large values.
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad$ $\forall t\in T$ (2i)
$\overline{c},\ \overline{d},\ \overline{e}\in\mathbb{R}_{\geq 0}\quad$ (3j)
@@
This model, however, contains many more constraints and variables than the others. A simplified investment and operation model is therefore also commonly used [22, 5]. We present tight formulations for simplified investment models in the online companion[20, Section 3].
@@ III Tight storage models
In this section, we explain the methodology and present the resulting tighter MIP formulations for the optimal investment and operation of storage problems (including reserves). In Section III-A, we first explain the outline of the process to obtain and prove the convex hull of the solutions to storage operation problems for one time period. We then apply this to the MIP models introduced in Section II to obtain tighter MIP models. The full proof for the first presented tight model can be found in Appendix -A. The proofs for the other tight models are in the online companion[20, Sections 5 and 6]. An overview of all models presented in this paper can also be found in the online companion[20, Section 1], as well as more tight MIP formulations for storage problems with different reserves and investments[20, Sections 2 – 4].
@@ III-A Method: Obtaining the Convex Hull
The convex hull of an MIP model is a set of constraints that form the tightest possible LP formulation of the solutions to this model. Every vertex of the convex hull is a feasible solution to the MIP model. For more background information on general LP and MIP theory, see [23]. In general, we cannot expect to explicitly generate the convex hull of an NP-hard problem, even if we allow for constraint types that contain exponentially many constraints [24]. However, for storage operation problems in one time period, we are able to find the convex hull by exploiting the disjunctive nature of the MIP. Balas[25] showed how the convex hull of any disjunctive problem can be obtained. This results in a convex hull formulation in a higher dimension than the original disjunctive problem, but it can be projected onto the original dimension. The obtained projected formulation is still a convex hull[23, Section 9.2.3]. Thus, it follows that the obtained formulation describes the convex hull of the original disjunctive problem.
@@
We can use the theory above to obtain the convex hull of storage problems in the following way. The first step is to write two disjunctive sets of constraints of the problem for one time period, where one describes the problem when charging, and the other when discharging. For the second step, we can then obtain the convex hull of these two disjunctive sets of constraints, as described Balas [25]. For a detailed explanation of how we did this, we refer to the proof in Appendix -A. However, the obtained convex hull formulation is in a higher dimension than the original formulation. Therefore, in the third step, we project the higher-dimensional formulation to the space of the original variables, by eliminating the extra variables using the Fourier-Motzkin elimination procedure [26, 27].
@@
The Fourier-Motzkin elimination procedure eliminates variables by combining all of their upper and lower bounds. The downside of this procedure is that it results in many extra constraints, of which some are facets (constraints describing the convex hull), but many others are redundant. It can be difficult to identify the redundant constraints and prove that they are indeed redundant, since all combinations of constraints need to be checked. However, it is typically beneficial to identify and remove redundant constraints, since this could reduce the size of the models. For example, for storage operation in one time period, we identified several redundant constraints in the convex hull found by Pozo[16]. As a result, we found a formulation for the same convex hull consisting of fewer constraints, as proven in Appendix -A.
@@
Many constraints are redundant when the parameters satisfy the following reasonable assumptions, which would typically hold in practice. It should always hold that $\overline{P^{\mathrm{C}}},R^{-}\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$ and $\overline{P^{\mathrm{D}}},R^{+}\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$, meaning that the charging/discharging capacity and the up/down reserve capacity for one time period cannot be larger than the total storage capacity of the storage unit. Obviously, $\overline{E}>\underline{E}\geq 0$ and $\overline{P^{\mathrm{C}}},\overline{P^{\mathrm{D}}},R^{+},R^{-}\geq 0$ here. For the efficiencies, $0<\eta^{\mathrm{C}},\eta^{\mathrm{D}}\leq 1$ should hold. Similarly, for the investment problem, it should always hold that $\overline{P^{\mathrm{C}}_{0}}+C\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}_{0}+E)(1-\phi)$ and $\overline{P^{\mathrm{D}}_{0}}+D\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}_{0}+E)(1-\theta)$, where $10\leq\phi<$ 1 and $\overline{P^{\mathrm{C}}_{0}},C,\overline{P^{\mathrm{D}}_{0}},D,\overline{E}_{0},E\geq 0$. Thus, to successfully replace some storage formulations with the tighter formulations presented in this paper, the parameters might need to be adapted to satisfy these requirements.
@@ III-B Tight MIP Formulation of Storage Operation
As a result of Theorem 1, the TO-MIP model is a tighter MIP model describing the optimal operation of storage than the original BO-MIP model. We illustrate this in Example 2. The TO-MIP model contains the facets (4b) and (4a), which were also found by Pozo [16]. However, the first advantage of the proposed TO-MIP model compared to the convex hull (in hyperplane representation) of Pozo is that the proposed model contains the same variables as the original MIP problem, whereas the variable $\delta_{t}$ is not present in the model by Pozo. Thus, the proposed formulation can be used to solve the original multi-period MIP storage operation problem. The second improvement is that we paid more attention to removing redundant constraints, which are not facets of the convex hull. As a result, the TO-LP model contains fewer constraints than their convex hull formulation. Note that if we apply Fourier-Motzkin in TO-MIP to eliminate variable $\delta_{t}$, then we can replace the upper bounds on $p^{\mathrm{C}}_{t}$ and $p^{\mathrm{D}}_{t}$ in (1c) and (1d) by $\frac{p^{\mathrm{C}}_{t}}{\overline{P^{\mathrm{C}}}}+\frac{p^{\mathrm{C}}_{t}}{\overline{P^{\mathrm{C}}}}\leq 1\ \forall t\in T$. This results in a model similar to that of Pozo, yet still smaller in size (4 constraints instead of 8).
@@ III-C Tight MIP Formulation of Storage Including Reserves
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad$ $\forall t\in T$ (2i)
@@
See the online companion[20, Section 5]. ∎
@@ III-D Tight MIP Formulation of Storage Investment Including Reserves
We now present the tighter models for storage investment and operation, including reserves. The formulations of these tighter models without reserves can simply be obtained by removing the reserve variables. For clarity, these tighter models are given in the online companion[20, Section 2]. We start by presenting the more general Tighter Investments and operation incl. Reserves MIP model (TIR-MIP model).
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad$ $\forall t\in T$ (2i)
$\overline{c},\ \overline{d},\ \overline{e}\in\mathbb{R}_{\geq 0}\quad$ (3j)
@@
See the online companion[20, Section 6]. ∎
@@ IV Illustrative case studies
To illustrate the improved performance of the newly obtained LP approximations, we apply these formulations to a unit commitment (UC) and a transmission expansion planning (TEP) case study, both presented by Arroyo et al. [6]. For a detailed description of the case studies, we refer to the paper by Arroyo [6]. We have slightly adapted these case studies such that they include reserves and investment decisions. Additionally, we have extended the two-period unit commitment case study to a multi-period problem. A detailed explanation of the adaptions to case studies is given in Appendix -B. We have implemented the case studies in Julia using JuMP [28], and solved them to optimality using the Gurobi solver [29, Version 10.0.2] with the default settings.
@@
In this section, we illustrate the performance of the proposed multi-period LP relaxation for storage operation (including reserves), and compare it to other LPs by embedding them in a unit commitment (UC) case study. More specifically, we compare the optimal solution of the basic MIP storage operation models BO-MIP and BOR-MIP, and of their LP relaxations, to the optimal solution of our LP relaxations TO-LP and TOR-LP. To this end, we replaced constraints (11)-(15) in the formulation by Wen et al. [10] in the case study by Arroyo et al. [6], which describe the basic operation of a storage unit, by the tighter storage operation formulations (with reserves) from this paper.
@@ IV-B LP Storage Investment in Transmission Expansion Planning Case Study
In this section, we illustrate the performance of our multi-period storage investment LP relaxation including reserves, and compare it to other LPs in a transmission expansion planning case study. More specifically, we compare the optimal solution of the basic MIP storage investment and operation model BIR-MIP, and of its LP relaxation, to the optimal solution of our LP relaxation TIR-LP. To this end, we replaced constraints (24)-(28) in the formulation presented by Arroyo et al. [6] (formulation originally by Zhou et al. [11]), which describe the basic operation of a storage unit, by the tighter storage investment and operation formulations (with reserves) from this paper.
@@ IV-C MIP Multi-Period Storage Operation in Unit Commitment Case Study
In this section, we illustrate the performance of the proposed multi-period MIP models for storage operation, and we compare them to basic MIP models by embedding them in a larger multi-period unit commitment case study. More specifically, we compare the computation time that the solver needs to solve the BO-MIP model to the time it needs for the TO-MIP model. To this end, we extend the original problem with two additional periods, in which the demand is 8.25MW and 0.0MW, respectively. We then repeat this problem 365 times (representing all days in one year), resulting in a problem consisting of 4$\cdot$365 = 1460 time periods. All other parameters remain the same. When solving this multi-period problem using an LP model for storage operation, the optimal solution favors simultaneous charging and discharging in many periods. This is for similar reasons as those discussed in Section IV-A, namely that charging and discharging simultaneously enables additional ramping, which is more cost-effective than starting up a second, very expensive generator. We now investigate how the models perform.
An overview of the results can be found in Table IV. It can be observed that the BO-MIP and TO-MIP models again find an optimal solution with the same objective function value, without simultaneous charging and discharging. More interestingly, the table displays the total time that is needed to solve the problem to optimality, averaged over 10 runs. Notice that less time is needed to solve the problem using the TO-MIP model. More concretely, fewer nodes are explored by the solver in the branch-and-cut algorithm for the TO-MIP model. This (partially) explains the difference in solve time. Thus, when solving a large-scale multi-period problem, the tighter TO-MIP model can indeed positively affect the solving time.
@@
The formulation by Pozo [16] cannot be used to solve this multi-period MIP problem, as their formulation does not contain the original binary variable. Instead, their model would yield results comparable to those of the TO-LP model, producing the same solution. However, the multi-period TO-LP formulation for this problem includes only half as many constraints as Pozo’s multi-period LP formulation.

### v3

@@ II-C Modeling Storage Investment Problems Including Reserves
We can extend the BOR-MIP model, as presented in the previous section, to model investment decisions. The most general way to model the investment of storage is the Basic Investment and operation incl. Reserves MIP model (BIR-MIP model) [morales-espanaImpactLargescaleHydrogen2024]. The operational problem is similar to that in the BOR-MIP model, but it contains additional investment variables $\overline{e}$, $\overline{c}$, and $\overline{d}$. The capacity limits of the storage unit, which now depend on the investment variables, are tracked in (3a) and (3b), where parameter $\overline{E}_{0}$ is the initially installed maximum storage capacity. Parameter $\theta$ represents the size of the minimum storage capacity as a fraction of the maximum. The upper bounds for charging and discharging and the reserves, also dependent on the investment variables, are imposed in (3c)-(3f), where parameters $\overline{P^{\mathrm{C}}_{0}}$ and $\overline{P^{\mathrm{D}}_{0}}$ are the initially installed charge and discharge capacities, and $C$ and $D$ are the maximum amount of invested charge/discharge capacity. Bounds on the maximum amount of invested capacity are imposed in (3g)-(3i), where $C$, $D$, and $E$ are typically quite large values.
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ $\forall t\in T$ (2i)
$\overline{c},\ \overline{d},\ \overline{e}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ (3j)
@@
This model, however, contains many more constraints and variables than the others. A simplified investment and operation model is therefore also commonly used [tejada-arangoPowerBasedGenerationExpansion2020, goAssessingEconomicValue2016]. We present tight formulations for simplified investment models in the online companion[elgersmaOnlineCompanionTight2024, Section 3].
@@ III Tight storage models
In this section, we explain the methodology and present the resulting tighter MIP formulations for the optimal investment and operation of storage problems (including reserves). In Section III-A, we first explain the outline of the process to obtain and prove the convex hull of the solutions to storage operation problems for one time period. We then apply this to the MIP models introduced in Section II to obtain tighter MIP models. The full proof for the first presented tight model can be found in Appendix -A. The proofs for the other tight models are in the online companion[elgersmaOnlineCompanionTight2024, Sections 5 and 6]. An overview of all models presented in this paper can also be found in the online companion[elgersmaOnlineCompanionTight2024, Section 1], as well as more tight MIP formulations for storage problems with different reserves and investments[elgersmaOnlineCompanionTight2024, Sections 2 – 4].
@@ III-A Method: Obtaining the Convex Hull
The convex hull of an MIP model is a set of constraints that form the tightest possible LP formulation of the solutions to this model. Every vertex of the convex hull is a feasible solution to the MIP model. For more background information on general LP and MIP theory, see [wolseyIntegerProgramming1998]. In general, we cannot expect to explicitly generate the convex hull of an NP-hard problem, even if we allow for constraint types that contain exponentially many constraints [papadimitriouComplexityFacetsFacets1982]. However, for storage operation problems in one time period, we are able to find the convex hull by exploiting the disjunctive nature of the MIP. Balas[balasDisjunctiveProgrammingHierarchy1985] showed how the convex hull of any disjunctive problem can be obtained. This results in a convex hull formulation in a higher dimension than the original disjunctive problem, but it can be projected onto the original dimension. The obtained projected formulation is still a convex hull[wolseyIntegerProgramming1998, Section 9.2.3]. Thus, it follows that the obtained formulation describes the convex hull of the original disjunctive problem.
@@
We can use the theory above to obtain the convex hull of storage problems in the following way. The first step is to write two disjunctive sets of constraints of the problem for one time period, where one describes the problem when charging, and the other when discharging. For the second step, we can then obtain the convex hull of these two disjunctive sets of constraints, as described Balas [balasDisjunctiveProgrammingHierarchy1985]. For a detailed explanation of how we did this, we refer to the proof in Appendix -A. However, the obtained convex hull formulation is in a higher dimension than the original formulation. Therefore, in the third step, we project the higher-dimensional formulation to the space of the original variables, by eliminating the extra variables using the Fourier-Motzkin elimination procedure [fourier1826solution, motzkin1936beitrage].
@@
The Fourier-Motzkin elimination procedure eliminates variables by combining all of their upper and lower bounds. The downside of this procedure is that it results in many extra constraints, of which some are facets (constraints describing the convex hull), but many others are redundant. It can be difficult to identify the redundant constraints and prove that they are indeed redundant, since all combinations of constraints need to be checked. However, it is typically beneficial to identify and remove redundant constraints, since this could reduce the size of the models. For example, for storage operation in one time period, we identified several redundant constraints in the convex hull found by Pozo[pozoConvexHullFormulations2023]. As a result, we found a formulation for the same convex hull consisting of fewer constraints, as proven in Appendix -A.
@@
Many constraints are redundant when the parameters satisfy the following reasonable assumptions, which would typically hold in practice. It should always hold that $\overline{P^{\mathrm{C}}},R^{-}\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}-\underline{E})$ and $\overline{P^{\mathrm{D}}},R^{+}\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}-\underline{E})$, meaning that the charging/discharging capacity and the up/down reserve capacity for one time period cannot be larger than the total storage capacity of the storage unit. Obviously, $\overline{E}>\underline{E}\geq 0$ and $\overline{P^{\mathrm{C}}},\overline{P^{\mathrm{D}}},R^{+},R^{-}\geq 0$ here. For the efficiencies, $0<\eta^{\mathrm{C}},\eta^{\mathrm{D}}\leq 1$ should hold. Similarly, for the investment problem, it should always hold that $\overline{P^{\mathrm{C}}_{0}}+C\leq\frac{1}{\eta^{\mathrm{C}}\Delta}(\overline{E}_{0}+E)(1-\phi)$ and $\overline{P^{\mathrm{D}}_{0}}+D\leq\frac{\eta^{\mathrm{D}}}{\Delta}(\overline{E}_{0}+E)(1-\theta)$, where $0\leq\phi<1$ and $\overline{P^{\mathrm{C}}_{0}},C,\overline{P^{\mathrm{D}}_{0}},D,\overline{E}_{0},E\geq 0$. Thus, to successfully replace some storage formulations with the tighter formulations presented in this paper, the parameters might need to be adapted to satisfy these requirements.
@@ III-B Tight MIP Formulation of Storage Operation
As a result of Theorem 1, the TO-MIP model is a tighter MIP model describing the optimal operation of storage than the original BO-MIP model. We illustrate this in Example 2. The TO-MIP model contains the facets (4b) and (4a), which were also found by Pozo [pozoConvexHullFormulations2023]. However, the first advantage of the proposed TO-MIP model compared to the convex hull (in hyperplane representation) of Pozo is that the proposed model contains the same variables as the original MIP problem, whereas the variable $\delta_{t}$ is not present in the model by Pozo. Thus, the proposed formulation can be used to solve the original multi-period MIP storage operation problem. The second improvement is that we paid more attention to removing redundant constraints, which are not facets of the convex hull. As a result, the TO-LP model contains fewer constraints than their convex hull formulation. Note that if we apply Fourier-Motzkin in TO-MIP to eliminate variable $\delta_{t}$, then we can replace the upper bounds on $p^{\mathrm{C}}_{t}$ and $p^{\mathrm{D}}_{t}$ in (1c) and (1d) by $\frac{p^{\mathrm{C}}_{t}}{\overline{P^{\mathrm{C}}}}+\frac{p^{\mathrm{C}}_{t}}{\overline{P^{\mathrm{C}}}}\leq 1\ \forall t\in T$. This results in a model similar to that of Pozo, yet still smaller in size (4 constraints instead of 8).
@@ III-C Tight MIP Formulation of Storage Including Reserves
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ $\forall t\in T$ (2i)
@@
See the online companion[elgersmaOnlineCompanionTight2024, Section 5]. ∎
@@ III-D Tight MIP Formulation of Storage Investment Including Reserves
We now present the tighter models for storage investment and operation, including reserves. The formulations of these tighter models without reserves can simply be obtained by removing the reserve variables. For clarity, these tighter models are given in the online companion[elgersmaOnlineCompanionTight2024, Section 2]. We start by presenting the more general Tighter Investments and operation incl. Reserves MIP model (TIR-MIP model).
@@
$e_{t},\ p^{\mathrm{C}}_{t},\ p^{\mathrm{D}}_{t},\ r^{\mathrm{C}+}_{t},\ r^{\mathrm{C}-}_{t},\ r^{\mathrm{D}+}_{t},\ r^{\mathrm{D}-}_{t}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ $\forall t\in T$ (2i)
$\overline{c},\ \overline{d},\ \overline{e}\in\mathbb{R}_{\geq 0}\quad\hskip-5.0pt$ (3j)
@@
See the online companion[elgersmaOnlineCompanionTight2024, Section 6]. ∎
@@ IV Illustrative case studies
To illustrate the improved performance of the newly obtained LP approximations, we apply these formulations to a unit commitment (UC) and a transmission expansion planning (TEP) case study, both presented by Arroyo et al. [arroyoUseConvexModel2020]. For a detailed description of the case studies, we refer to the paper by Arroyo [arroyoUseConvexModel2020]. We have slightly adapted these case studies such that they include reserves and investment decisions. Additionally, we have extended the two-period unit commitment case study to a multi-period problem. A detailed explanation of the adaptions to case studies is given in Appendix -B. We have implemented the case studies in Julia using JuMP [Lubin2023], and solved them to optimality using the Gurobi solver [gurobi, Version 10.0.2] with the default settings.
@@
In this section, we illustrate the performance of the proposed multi-period LP relaxation for storage operation (including reserves), and compare it to other LPs by embedding them in a unit commitment (UC) case study. More specifically, we compare the optimal solution of the basic MIP storage operation models BO-MIP and BOR-MIP, and of their LP relaxations, to the optimal solution of our LP relaxations TO-LP and TOR-LP. To this end, we replaced constraints (11)-(15) in the formulation by Wen et al. [wenEnhancedSecurityConstrainedUnit2016] in the case study by Arroyo et al. [arroyoUseConvexModel2020], which describe the basic operation of a storage unit, by the tighter storage operation formulations (with reserves) from this paper.
@@ IV-B LP Storage Investment in Transmission Expansion Planning Case Study
In this section, we illustrate the performance of our multi-period storage investment LP relaxation including reserves, and compare it to other LPs in a transmission expansion planning case study. More specifically, we compare the optimal solution of the basic MIP storage investment and operation model BIR-MIP, and of its LP relaxation, to the optimal solution of our LP relaxation TIR-LP. To this end, we replaced constraints (24)-(28) in the formulation presented by Arroyo et al. [arroyoUseConvexModel2020] (formulation originally by Zhou et al. [zhaoUsingElectricalEnergy2018]), which describe the basic operation of a storage unit, by the tighter storage investment and operation formulations (with reserves) from this paper.
@@ IV-C MIP Multi-Period Storage Operation in Unit Commitment Case Study
[…]
An overview of the results can be found in Table IV. It can be observed that the BO-MIP and TO-MIP models again find an optimal solution with the same objective function value, without simultaneous charging and discharging. More interestingly, the table displays the total time that is needed to solve the problem to optimality, averaged over 10 runs. Notice that less time is needed to solve the problem using the TO-MIP model. More concretely, fewer nodes are explored by the solver in the branch-and-cut algorithm for the TO-MIP model. This (partially) explains the difference in solve time. Thus, when solving a large-scale multi-period problem, the tighter TO-MIP model could indeed positively affect the solving time.
@@
The formulation by Pozo [pozoConvexHullFormulations2023] cannot be used to solve this multi-period MIP problem, as their formulation does not contain the original binary variable. Instead, their model would yield results comparable to those of the TO-LP model, producing the same solution. However, the multi-period TO-LP formulation for this problem includes only half as many constraints as Pozo’s multi-period LP formulation.